Find the inverse of the matrix (if it exists). (If an answer does not exist, enter DNE.)
Answer #1

1 2 5 9 let A= a 2 A-E JE 5 9 a=1 b !) 2 C=5 d=9 - 1 Al=ad-bc = (txt) - x)= Al=-1 to son singulas mateix Inverse matrix Inver